# Large Language Models Miss the Multi-Agent Mark

### v3

@@ 1 Introduction
The recent machine learning literature has seen an upsurge in popularity of Large Language Models (LLMs) used in coordination to solve complex tasks, a line of research that goes by the name of “Multi-Agent Systems of LLMs” (MAS LLMs) (guo2024largelanguagemodel, ; tran2025multiagentcollaborationmechanismssurvey, ). In MAS LLMs, each LLM-agent specialises in a task to accomplish a goal. A few examples of MAS LLMs use are software engineering (holt2024l2maclargelanguagemodel, ), multi-robot planning (chen2024scalable, ), data analysis (rasheed2024largelanguagemodelsserve, ), scientific production, reasoning and debating (du2023improvingfactualityreasoninglanguage, ; tang2024medagentslargelanguagemodels, ; zheng2023chatgptresearch, ), and social simulations (park2023generative, ), among many others. There is also an increasing interest in open-ended MAS LLMs, systems whose complex interactions give rise to human-like emergent behaviours (al2024projectsidmanyagentsimulations, ; gao2023s3socialnetworksimulationlarge, ; park2023generative, ).
However, labelling MAS LLMs as “Multi-Agent Systems” has already raised concerns in the scientific community (chen2024llmcallsneedscaling, ). Influential frameworks employed in MAS LLMs applications, such as ReAct (yao2023reactsynergizingreasoningacting, ), which in turn leverages methods like Chain of Thought (wei2023chainofthoughtpromptingelicitsreasoning, ), Tree of Thought (yao2023treethoughtsdeliberateproblem, ), etc., are single-agent prompting techniques that overlook concurrency and shared states; [^1] https://gist.github.com/yoavg/9142e5d974ab916462e8ec080407365b At the same time, agentic frameworks developed by large companies are monolithic orchestrators that leverage (and only cite) machine learning research, taking little notice of decades of MAS research.2
While the term MAS LLMs was introduced in 2023 (talebirad2023multi, ), the first MAS works date back to the 1980s and the 1990s (10.1109/69.43404, ; shoham1993agent, ; wooldridge1995intelligentagents, ). In this sense, our position advocates for using precise scientific terminology and cautions against the risk of reinventing the wheel. For a relatively new field like that of MAS LLMs, failing to engage with the broader MAS literature may lead to overlooked insights and missed opportunities for meaningful advancement.
@@
I. Social intelligent agents: LLM agents lack native social behaviour. MAS agents populate an environment and receive high-level goals to fulfil. Such goals necessitate the realisation, specification, and completion of other possibly unanticipated sub-tasks. In this context, a high-level goal tests an agent’s reactivity, proactiveness, and social abilities (wooldridge1995intelligentagents, ). A reactive agent dynamically perceives the environment and takes the initiative to satisfy its design objectives. In doing so, a social agent is capable of interacting and intelligently competing with other agents.
In the context of LLMs, agents are both reactive and proactive: they exhibit remarkable adaptability to changes of an input prompt (which account for their environment, in most cases), and can be trained to take the initiative on how to split a task into its elementary components and then complete them, as shown by frameworks such as AutoGPT (yang2023autogptonlinedecisionmaking, ).
In the current MAS LLM literature, we highlight the lack of consideration given to the social aspects of reactivity and proactiveness. While collaborating and competing are prerequisites of any MAS, most LLMs are fine-tuned as single agents, with no proper multi-agent pre-training procedure (kim2023multipromptercooperativepromptoptimization, ). [^2] https://www.kaggle.com/whitepaper-agents, https://www.anthropic.com/engineering/building-effective-agents, https://openai.com/index/new-tools-for-building-agents/ In other words, LLMs are trained in isolation to respond to users’ requests, rather than to interact with each other. This leads to poor performance and unexpected failures when LLMs are benchmarked to identify other agents’ beliefs, desires and intentions, i.e., with Theory of Mind problems (shapira2023cleverhansneuraltheory, ; ullman2023largelanguagemodelsfail, ).
These simplifications limit the field’s potential and risk misdirecting current efforts toward MAS, when in some cases the problem may be better addressed by aggregation methods that combine many independent components, such as ensembles (du2024multiagentsoftwaredevelopmentcrossteam, ; kapoor2024aiagentsmatter, ; li2024agentsneed, ; su2025headsbetteroneimproved, ). [^3] The authors of “More agents is all you need” make explicit that their technique is an ensemble, as per their Figure 1 (li2024agentsneed, ).
@@
II. Environment design: MAS LLMs environments are LLM-centric. MAS traditionally model the environment with no strong assumption on the architecture or configuration of the agents that populate it (cossentino2005requirements, ; park2005designing, ; wooldridge2000gaia, ). Conversely, MAS LLMs subvert this approach: the environment design assumes the agents are LLMs that communicate and coordinate via natural language. This Section warns that this assumption overlooks the intrinsic limitations of LLMs and argues for addressing such problems by designing MAS environments that are not LLM-centric.
Consider the environment predictability:; thus, a fully deterministic environment cannot exist when involving LLMs, providing little control over procedures one wants to guarantee to be safe or terminate in a specific state.
Further,. To foster effective cooperation and competition, that should not reduce to the environment but include representing other agents’ actions, beliefs and intentions. Unfortunately, LLMs are not only known to struggle with inferring beliefs and intentions (sclar2023mindinglanguagemodelslack, ; ullman2023largelanguagemodelsfail, ), a long-standing challenge in the MAS literature (10.1007/978-3-031-45368-7_24, ; shi2025mumatommultimodalmultiagenttheory, ), but also with maintaining a consistent representation of the environment due to issues like hallucinations and memory persistency.
Another critical aspect of MAS is how the environment is represented and then perceived by the agents. In most MAS LLMs settings, the environment is textual or translated as such, with text being the medium of reference for the largest majority of open- and closed-source LLMs. Storing such representations may rapidly exceed a model’s context length and cause hallucinations (li2023camel, ; li2023metaagentssimulatinginteractionshuman, ).
To realise LLMs as MAS agents that work in coordination with humans and other non-LLM agents, we should design open, multi-modal environments that address their lack of long-term memory persistency (zhang2024surveymemorymechanismlarge, ), non-determinism and propensity to hallucinations (lamalfa2023jair, ), and the costs and intrinsic ambiguity of natural language as a storage and communication medium (marro2024scalablecommunicationprotocolnetworks, ).
@@
Asynchronicity is typically absent in MAS LLMs, as LLM-agents often operate in strictly sequential pipelines or parallel, rather than as independent, concurrently operating agents. While some works are moving in this direction (ginart2024asynchronoustoolusagerealtime, ), we argue for leveraging the body of knowledge from the field of multi-agent concurrent systems that is operative since the 1990s (see, for example, (135681, ; 10.1007/3-540-58855-8_1, )).
Another fundamental aspect of MAS and MAS LLMs is communication between agents. Most MAS LLMs literature assumes agents communicate via natural language (qin2024largelanguagemodelsmeet, ). This smoothness is a simplification that overlooks the complexity of real agent interaction and decades of research in MAS communication systems and protocols (berna2004communication, ; jang2005efficient, ; saunders1996evolution, ).
Natural language communication is, in fact, inefficient and ambiguous: standards that describe agents’ beliefs, commitments, and actions. In line with the MAS literature, we argue mechanisms to negotiate and implement communication methods that integrate the principles of speech acts (austin1962how, ; shoambook, ) and Gricean maxims (Grice1975-GRILAC-6, ) to minimise the cost of communication and maximise its effectiveness.
@@
IV. MAS LLMs do not quantify the emergence of complex behaviours. Several recent works considered MAS LLMs as open-ended systems (al2024projectsidmanyagentsimulations, ; park2023generative, ; xie2024largelanguagemodelagents, ), i.e., scenarios where the long-term evolution of the system produces complex behaviours and solutions (openendedlearningteam2021openendedlearningleadsgenerally, ; wang2019pairedopenendedtrailblazerpoet, ). Intuitively, large-scale, long-term interactions would play a similar role as size (the so-called “scaling laws for LLMs” (kaplan2020scalinglawsneurallanguage, )) in shaping complex behaviour and dynamics that would not emerge otherwise.
Currently, despite growing hype around the potential of LLMs in MAS contexts, which often exceed their demonstrated abilities (shanahan2023talkinglargelanguagemodels, ), we foresee the risk of a birth and death of interest in emergent behaviours in LLMs. Emergence alone, primarily when arising from loosely constrained prompts and undefined interaction dynamics, further compromised by hallucinations and memory issues in LLMs, does not justify claims of MAS-level coordination or long-horizon planning. These factors make distinguishing between genuine coordination and coincidental or spurious outputs increasingly difficult.
We therefore argue that the evaluation of emergent behaviours in LLMs should rely on quantifiable metrics rooted in established MAS (balduzzi2019openendedlearningsymmetriczerosum, ; NEURIPS2023_764c18ad, ; 10.5555/3306127.3331691, ; NEURIPS2022_b2a1c152, ).
@@ 2 Social Intelligent Agents: LLM Agents Lack Native Social Behaviour
In MAS, an agent is commonly defined as intelligent if it is reactive, proactive, and shows social behaviours (russel2010, ; wooldridge, ). While LLMs are reactive and proactive, this section discusses why they often fail to be social agents and where this discrepancy originates.
@@
For any agent, the prerequisite of reactiveness is perception, i.e., interpreting an input to take an associated action. LLMs are reactive agents, i.e., respond to stimuli from a changing environment, with the input being directly injected into the prompt (shankar2024buildingreactive, ; sinha2024realtimeanomalydetectionreactive, ; yang2024text2reaction, ) or retrieved via external information (jin2024reasoninggraspingmultimodallarge, ; sun2024interactiveplanning, ).
@@
The proactiveness of an LLM is its capacity to initiate tasks with limited or without human intervention (lu2024proactiveagentshiftingllm, ). Independence eventually arises in LLMs only when the model can self-generate the prompt or self-inject information (ouyang2022traininglanguagemodelsfollow, ). Recent works have focused on proactive LLMs that ask clarification to the user on the instructions provided, retrieve additional information from external resources or are fine-tuned to anticipate the next actions (zeighami2025llmpoweredproactivedatasystems, ; zhang2024askbeforeplanproactivelanguageagents, ).
@@
The third intelligent characteristic of an agent is the capacity to behave socially, thus to compete or cooperate (conitzer2023foundationsof, ; wooldridge1995intelligentagents, ). Competition and cooperation arise from the capacity to be both reactive and proactive, i.e., react to other agents’ actions and initiate new ones. Crucially, agents must be socially reactive and proactive, i.e., able to grasp and reason about other agents’ goals, negotiate with them, and even enlist their cooperation when needed.
@@
The literature on LLMs includes a substantial body. In most of them, the agents’ roles are scripted through prompting (du2023improvingfactualityreasoninglanguage, ; mosquera2024llmaugmentedautonomousagentscooperate, ) or fine-tuned (ma2025coevolvingyoufinetuningllm, ; subramaniam2025multiagentfinetuningselfimprovement, ; yang2024multiobjectivefinetuningenhancedprogram, ), but not natively trained to cooperate or compete with one another.
@@
A recent survey, identified that $37\%$ of the failure cases of MAS LLMs are errors caused by inter-agents misalignment or agents’ coordination issues (cemri2025multiagentllmsystemsfail, ) (Figure 2). In the same spirit, (Li_2023, ) shows how multi-agent reinforcement learning agents outperform LLMs at planning tasks, highlighting the role of pre-training. Concurrently, a growing body of research in Machine Theory of Mind (ToM) (rabinowitz2018machinetheorymind, ) evidences how LLMs struggle to express their beliefs, desires, and intentions (d2004dmars, ; rao1995bdi, ), and those of other agents (strachan2024testing, ; ullman2023largelanguagemodelsfail, ; van2023theory, ).
Furthermore, when multiple LLMs interact,.
@@
Recent advances in the field allow training LLMs directly on textual feedback they receive from other models (yuksekgonul2024textgradautomaticdifferentiationtext, ); that, alongside cooperative or competitive reinforcement learning, where agents are trained not only to complete tasks but also to adapt to and influence each other’s behaviour, can serve as a path towards MAS in which LLMs show reactive and proactive behaviours beyond answering to prompts. An agent trained in this way would learn to respond appropriately while also interpreting, anticipating, and reacting to the actions of other agents, based on feedback from their interactions.
In conclusion, while fine-tuning is proven to be promising to specialise and assign roles to LLMs (liao2025marftmultiagentreinforcementfinetuning, ; ma2025coevolvingyoufinetuningllm, ; subramaniam2025multiagentfinetuningselfimprovement, ), we argue it is alone insufficient to provide them with the capacity to cooperate and compete.

### v4

@@ 1 Introduction
The recent machine learning literature has seen an upsurge in popularity of Large Language Models (LLMs) used in coordination to solve complex tasks, a line of research that goes by the name of “Multi-Agent Systems of LLMs” (MAS LLMs) (guo2024largelanguagemodel; tran2025multiagentcollaborationmechanismssurvey). In MAS LLMs, each LLM-agent specialises in a task to accomplish a goal. A few examples of MAS LLMs use are software engineering (holt2024l2maclargelanguagemodel), multi-robot planning (chen2024scalable), data analysis (rasheed2024largelanguagemodelsserve), scientific production, reasoning and debating (du2023improvingfactualityreasoninglanguage; tang2024medagentslargelanguagemodels; zheng2023chatgptresearch), and social simulations (park2023generative), among many others. There is also an increasing interest in open-ended MAS LLMs, systems whose complex interactions give rise to human-like emergent behaviours (al2024projectsidmanyagentsimulations; gao2023s3socialnetworksimulationlarge; park2023generative).
However, labelling MAS LLMs as “Multi-Agent Systems” has already raised concerns in the scientific community (chen2024llmcallsneedscaling). Influential frameworks employed in MAS LLMs applications, such as ReAct (yao2023reactsynergizingreasoningacting), which in turn leverages methods like Chain of Thought (wei2023chainofthoughtpromptingelicitsreasoning), Tree of Thought (yao2023treethoughtsdeliberateproblem), etc., are single-agent prompting techniques that overlook concurrency and shared states; [^1] https://gist.github.com/yoavg/9142e5d974ab916462e8ec080407365b At the same time, agentic frameworks developed by large companies are monolithic orchestrators that leverage (and only cite) machine learning research, taking little notice of decades of MAS research.2
While the term MAS LLMs was introduced in 2023 (talebirad2023multi), the first MAS works date back to the 1980s and the 1990s (10.1109/69.43404; shoham1993agent; wooldridge1995intelligentagents). In this sense, our position advocates for using precise scientific terminology and cautions against the risk of reinventing the wheel. For a relatively new field like that of MAS LLMs, failing to engage with the broader MAS literature may lead to overlooked insights and missed opportunities for meaningful advancement.
@@
I. Social intelligent agents: LLM agents lack native social behaviour. MAS agents populate an environment and receive high-level goals to fulfil. Such goals necessitate the realisation, specification, and completion of other possibly unanticipated sub-tasks. In this context, a high-level goal tests an agent’s reactivity, proactiveness, and social abilities (wooldridge1995intelligentagents). A reactive agent dynamically perceives the environment and takes the initiative to satisfy its design objectives. In doing so, a social agent is capable of interacting and intelligently competing with other agents.
In the context of LLMs, agents are both reactive and proactive: they exhibit remarkable adaptability to changes of an input prompt (which account for their environment, in most cases), and can be trained to take the initiative on how to split a task into its elementary components and then complete them, as shown by frameworks such as AutoGPT (yang2023autogptonlinedecisionmaking).
In the current MAS LLM literature, we highlight the lack of consideration given to the social aspects of reactivity and proactiveness. While collaborating and competing are prerequisites of any MAS, most LLMs are fine-tuned as single agents, with no proper multi-agent pre-training procedure (kim2023multipromptercooperativepromptoptimization). [^2] https://www.kaggle.com/whitepaper-agents, https://www.anthropic.com/engineering/building-effective-agents, https://openai.com/index/new-tools-for-building-agents/ In other words, LLMs are trained in isolation to respond to users’ requests, rather than to interact with each other. This leads to poor performance and unexpected failures when LLMs are benchmarked to identify other agents’ beliefs, desires and intentions, i.e., with Theory of Mind problems (shapira2023cleverhansneuraltheory; ullman2023largelanguagemodelsfail).
These simplifications limit the field’s potential and risk misdirecting current efforts toward MAS, when in some cases the problem may be better addressed by aggregation methods that combine many independent components, such as ensembles (du2024multiagentsoftwaredevelopmentcrossteam; kapoor2024aiagentsmatter; li2024agentsneed; su2025headsbetteroneimproved). [^3] The authors of “More agents is all you need” make explicit that their technique is an ensemble, as per their Figure 1 (li2024agentsneed).
@@
II. Environment design: MAS LLMs environments are LLM-centric. MAS traditionally model the environment with no strong assumption on the architecture or configuration of the agents that populate it (cossentino2005requirements; park2005designing; wooldridge2000gaia). Conversely, MAS LLMs subvert this approach: the environment design assumes the agents are LLMs that communicate and coordinate via natural language. This Section warns that this assumption overlooks the intrinsic limitations of LLMs and argues for addressing such problems by designing MAS environments that are not LLM-centric.
Consider the environment predictability: for an environment that is meant to be deterministic, LLMs are inherently not (lamalfa2023jair; Ouyang_2025); thus, a fully deterministic environment cannot exist when involving LLMs, providing little control over procedures one wants to guarantee to be safe or terminate in a specific state.
Further, agents should receive and consistently maintain a unique representation of the environment (Durfee2001; rao1995bat), particularly when settings are dynamic and partially observable (10.5555/2967142). To foster effective cooperation and competition, that should not reduce to the environment but include representing other agents’ actions, beliefs and intentions. Unfortunately, LLMs are not only known to struggle with inferring beliefs and intentions (sclar2023mindinglanguagemodelslack; ullman2023largelanguagemodelsfail), a long-standing challenge in the MAS literature (10.1007/978-3-031-45368-7_24; shi2025mumatommultimodalmultiagenttheory), but also with maintaining a consistent representation of the environment due to issues like hallucinations and memory persistency.
Another critical aspect of MAS is how the environment is represented and then perceived by the agents. In most MAS LLMs settings, the environment is textual or translated as such, with text being the medium of reference for the largest majority of open- and closed-source LLMs. Storing such representations may rapidly exceed a model’s context length and cause hallucinations (li2023camel; li2023metaagentssimulatinginteractionshuman).
To realise LLMs as MAS agents that work in coordination with humans and other non-LLM agents, we should design open, multi-modal environments that address their lack of long-term memory persistency (zhang2024surveymemorymechanismlarge), non-determinism and propensity to hallucinations (lamalfa2023jair), and the costs and intrinsic ambiguity of natural language as a storage and communication medium (marro2024scalablecommunicationprotocolnetworks).
@@
Asynchronicity is typically absent in MAS LLMs, as LLM-agents often operate in strictly sequential pipelines or parallel, rather than as independent, concurrently operating agents. While some works are moving in this direction (ginart2024asynchronoustoolusagerealtime), we argue for leveraging the body of knowledge from the field of multi-agent concurrent systems that is operative since the 1990s (see, for example, (135681; 10.1007/3-540-58855-8_1)).
Another fundamental aspect of MAS and MAS LLMs is communication between agents. Most MAS LLMs literature assumes agents communicate via natural language (qin2024largelanguagemodelsmeet). This smoothness is a simplification that overlooks the complexity of real agent interaction and decades of research in MAS communication systems and protocols (berna2004communication; jang2005efficient; saunders1996evolution).
Natural language communication is, in fact, inefficient and ambiguous: in contrast, MAS communication protocols in the form of structured languages (e.g., KQML (kqml1994) or Agent-Oriented Programming (shoham1993agent)) are consolidated performative standards that describe agents’ beliefs, commitments, and actions. In line with the MAS literature, we argue mechanisms to negotiate and implement communication methods that integrate the principles of speech acts (austin1962how; shoambook) and Gricean maxims (Grice1975-GRILAC-6) to minimise the cost of communication and maximise its effectiveness.
@@
IV. MAS LLMs do not quantify the emergence of complex behaviours. Several recent works considered MAS LLMs as open-ended systems (al2024projectsidmanyagentsimulations; park2023generative; xie2024largelanguagemodelagents), i.e., scenarios where the long-term evolution of the system produces complex behaviours and solutions (openendedlearningteam2021openendedlearningleadsgenerally; wang2019pairedopenendedtrailblazerpoet). Intuitively, large-scale, long-term interactions would play a similar role as size (the so-called “scaling laws for LLMs” (kaplan2020scalinglawsneurallanguage)) in shaping complex behaviour and dynamics that would not emerge otherwise.
Currently, despite growing hype around the potential of LLMs in MAS contexts, which often exceed their demonstrated abilities (shanahan2023talkinglargelanguagemodels), we foresee the risk of a birth and death of interest in emergent behaviours in LLMs. Emergence alone, primarily when arising from loosely constrained prompts and undefined interaction dynamics, further compromised by hallucinations and memory issues in LLMs, does not justify claims of MAS-level coordination or long-horizon planning. These factors make distinguishing between genuine coordination and coincidental or spurious outputs increasingly difficult.
We therefore argue that the evaluation of emergent behaviours in LLMs should rely on quantifiable metrics rooted in established MAS (balduzzi2019openendedlearningsymmetriczerosum; NEURIPS2023_764c18ad; 10.5555/3306127.3331691; NEURIPS2022_b2a1c152).
@@ 2 Social Intelligent Agents: LLM Agents Lack Native Social Behaviour
In MAS, an agent is commonly defined as intelligent if it is reactive, proactive, and shows social behaviours (russel2010; wooldridge). While LLMs are reactive and proactive, this section discusses why they often fail to be social agents and where this discrepancy originates.
@@
For any agent, the prerequisite of reactiveness is perception, i.e., interpreting an input to take an associated action. LLMs are reactive agents, i.e., respond to stimuli from a changing environment, with the input being directly injected into the prompt (shankar2024buildingreactive; sinha2024realtimeanomalydetectionreactive; yang2024text2reaction) or retrieved via external information (jin2024reasoninggraspingmultimodallarge; sun2024interactiveplanning).
@@
The proactiveness of an LLM is its capacity to initiate tasks with limited or without human intervention (lu2024proactiveagentshiftingllm). Independence eventually arises in LLMs only when the model can self-generate the prompt or self-inject information (ouyang2022traininglanguagemodelsfollow). Recent works have focused on proactive LLMs that ask clarification to the user on the instructions provided, retrieve additional information from external resources or are fine-tuned to anticipate the next actions (zeighami2025llmpoweredproactivedatasystems; zhang2024askbeforeplanproactivelanguageagents).
@@
The third intelligent characteristic of an agent is the capacity to behave socially, thus to compete or cooperate (conitzer2023foundationsof; wooldridge1995intelligentagents). Competition and cooperation arise from the capacity to be both reactive and proactive, i.e., react to other agents’ actions and initiate new ones. Crucially, agents must be socially reactive and proactive, i.e., able to grasp and reason about other agents’ goals, negotiate with them, and even enlist their cooperation when needed.
@@
The literature on LLMs includes a substantial body of works aimed at developing competitive and cooperative systems of LLMs (liu2023bolaabenchmarkingorchestratingllmaugmented). In most of them, the agents’ roles are scripted through prompting (du2023improvingfactualityreasoninglanguage; mosquera2024llmaugmentedautonomousagentscooperate) or fine-tuned (ma2025coevolvingyoufinetuningllm; subramaniam2025multiagentfinetuningselfimprovement; yang2024multiobjectivefinetuningenhancedprogram), but not natively trained to cooperate or compete with one another.
@@
A recent survey identified that $37\%$ of the failure cases of MAS LLMs are errors caused by inter-agents misalignment or agents’ coordination issues (cemri2025multiagentllmsystemsfail) (Figure 2). In the same spirit, (Li_2023) shows how multi-agent reinforcement learning agents outperform LLMs at planning tasks, highlighting the role of pre-training. Concurrently, a growing body of research in Machine Theory of Mind (ToM) (rabinowitz2018machinetheorymind) evidences how LLMs struggle to express their beliefs, desires, and intentions (d2004dmars; rao1995bdi), and those of other agents (strachan2024testing; ullman2023largelanguagemodelsfail; van2023theory).
Furthermore, when multiple LLMs interact, their specialisation, achieved through different prompts or fine-tuning, often aims to maximise a single objective such as accuracy, at the expense of other desirable behavioural properties. As a consequence, several MAS LLMs reduce to aggregation mechanisms like majority vote (du2024multiagentsoftwaredevelopmentcrossteam; kapoor2024aiagentsmatter; li2024agentsneed; su2025headsbetteroneimproved).
@@
Recent advances in the field allow training LLMs directly on textual feedback they receive from other models (yuksekgonul2024textgradautomaticdifferentiationtext); that, alongside cooperative or competitive reinforcement learning, where agents are trained not only to complete tasks but also to adapt to and influence each other’s behaviour, can serve as a path towards MAS in which LLMs show reactive and proactive behaviours beyond answering to prompts. An agent trained in this way would learn to respond appropriately while also interpreting, anticipating, and reacting to the actions of other agents, based on feedback from their interactions.
In conclusion, while fine-tuning is proven to be promising to specialise and assign roles to LLMs (liao2025marftmultiagentreinforcementfinetuning; ma2025coevolvingyoufinetuningllm; subramaniam2025multiagentfinetuningselfimprovement), we argue it is alone insufficient to provide them with the capacity to cooperate and compete.